OpenStudy (anonymous):

-2+square root of 20 over 2

OpenStudy (anonymous):

\[\frac{-2+\sqrt{20}}{2}\]

OpenStudy (anonymous):

break it up. \[\frac{-2}{2}+\frac{\sqrt{20}}{2} = -1+\frac{\sqrt{4*5}}{2} = -1 + \frac{\sqrt{2^{2}*5}}{2}=-1+\frac{2\sqrt{5}}{2} =1+\sqrt{5}\]